# Probing Saturon-like Limits in QCD Systems

### v1

@@ III.3 Discussions
Based on the results presented above and under the assumptions considered, while both the formal structure and the intrinsic nature of nonlinear QCD evolution at small $x$ appear to support the prospect of the proton as a candidate saturon system, current evaluations–whether from the perspective of gluon occupation number or thermodynamic entropy–indicate that the criteria required to reach the upper bound for saturon identification remain stringent. This necessitates collisions at higher energies and gluonic states with greater occupation densities. In contrast, once the same analysis chain is applied to nuclear targets (fixed coupling, identical numerical settings otherwise), the nuclear curves rise faster. The resulting nuclear entropy $S_{A}(x)$ approaches and, within the uncertainties of our inputs, can reach the benchmark line $1/\alpha_{s}$ in a small-$x$ window where the proton stays below. This qualitative separation is stable against reasonable variations of the $Q_{s}$ extraction threshold, geometric inputs, and numerical smoothing; these choices primarily shift the apparent crossing point but do not invert the ordering. Encouragingly, some studies have already explored the correspondence between black hole saturons and the CGC framework [12]. Portions of those discussions align conceptually with the present work, and notably, they propose a role for Goldstone modes arising from the breaking of Poincaré symmetry. The breaking of Poincaré invariance in this context fundamentally originates from the presence of a hard scale, which characterizes the transition from a dilute partonic gas with a spectral gap to an over-occupied, highly saturated classical field. Partons with momenta below this scale become screened, and high-twist effects become significant. This screening of partonic modes disrupts the translational invariance of the partonic gas, analogous to the breaking of Poincaré symmetry in black holes, which gives rise to a Goldstone scale characterized by $f=m_{\text{Planck}}$ with Planck mass $m_{\text{Planck}}$ [3]. However, in the case of QCD, the precise form and interpretation of such a Goldstone scale remain open to further investigation.
@@
In this work, we explored whether a high-energy proton in the small-$x$ regime could qualify as a saturon-a highly occupied field configuration that maximizes entropy under unitarity constraints, in analogy with black holes. We derived the gluon occupation number $N_{g}(x)$ and a thermodynamic entropy $S(x)$ based on analytical and numerical solutions to the BK equation (with both fixed and running coupling), supplemented with an Unruh-inspired temperature $T=Q_{s}/(2\pi)$ and an emergent gluon mass $M_{g}\sim Q_{s}$. Our findings indicate that although both $N_{g}$ and $S$ increase as $x$ decreases, even down to $x\sim 10^{-7}$, they remain significantly below the theoretical saturon bound $1/\alpha_{s}$. This suggests that while the proton exhibits nonlinear saturation dynamics, under the current theoretical framework it does not reach the critical maximal occupancy characteristic of a saturon. After extending the same framework to nuclear targets at fixed coupling, the nuclear entropy $S_{A}(x)$ rises more sharply and, in our numerical study, can reach the $1/\alpha_{s}$ benchmark in a region where the proton does not. This proton–nucleus contrast identifies nuclei as more favorable environments to probe saturon-like behavior in QCD. Moreover, exploring the connection between the proton and black holes from the perspective of entanglement entropy offers a promising avenue [13]. Notably, some studies have already applied the entanglement-based definition originally proposed for black holes to DIS processes involving protons.
To advance toward the saturon regime, we propose two complementary pathways. On the theoretical front, it is essential to incorporate higher-order QCD corrections-such as NLO effects or small-$x$ resummations-and to explore alternative entropy constructs, particularly entanglement entropy, which may more accurately capture the gluonic microstate structure. Investigating symmetry-breaking phenomena in the saturated regime, such as emergent Goldstone-like modes, may also uncover universal behavior reminiscent of saturon systems.
From experimental and phenomenological perspective, heavy-ion collisions offer a promising environment to probe truly over-occupied gluonic systems. The CGC effective theory, which describes saturated gluon matter at high energy, has been successfully applied to model the initial stages of heavy-ion collisions at RHIC and the LHC [88]. These systems, in which gluon densities are far higher than in isolated protons, may approach saturon-like occupancy. Additionally, future facilities such as the Electron-Ion Collider (EIC) are expected to explore deep small-$x$ regimes in both proton and nuclear targets, with unprecedented luminosity and kinematic reach [89, 90, 91, 92]. A comprehensive study-spanning proton-proton, proton-nucleus, and nucleus-nucleus collisions-of gluon saturation, entropy production, and correlation patterns could shed light on whether any QCD system realizes saturon-like behavior. Such investigations might reveal deep connections between hadronic saturation dynamics and gravitational entropy bounds.
@@ Acknowledgments
This work has been supported by the National Key R&\&D Program of China (Grant NO. 2024YFE0109800 and 2024YFE0109802).

### v2

@@ III.3 Discussions
Based on the results presented above, the proton appears to fall short of the saturon bound ($S\sim 1/\alpha_{s}$) within the considered kinematic range. We identify three key physical factors contributing to this observation:
1. Geometric Dilution and Finite-Size Evolution: Our measure of the average occupancy (Eq. 12) involves an integration over the transverse area. Since the proton has a non-uniform density profile, the dilute periphery lowers the average. Moreover, as recently characterized in Ref. [95], the geometry of the target is dynamic. The evolution equation with impact-parameter dependence captures finite-size effects (Gribov diffusion), which leads to a spatial expansion of the proton. This study demonstrates that neglecting these finite-size effects overestimates saturation in protons. Therefore, the dynamical evolution of the geometry acts as an additional mechanism that suppresses the growth of entropy density in the proton, reinforcing our conclusion that the saturon limit is difficult to reach in e-p collisions compared to heavy nuclei.
2. Suppression from Higher-Order Effects: The Leading Order (LO) approximation used in standard BK evolution often overestimates the growth speed. It is established that including Next-to-Leading Log (NLL) corrections and small-$x$ resummations significantly reduces the intercept of the gluon evolution [96, 97]. Consequently, in a full higher-order calculation, the entropy would likely grow even more slowly than in our current analysis, reinforcing the conclusion that reaching the saturon limit is extremely challenging for a single proton.
3. Entanglement vs. Thermodynamics: The thermodynamic entropy calculated here (Eq. 18) is a semiclassical proxy. A full quantum mechanical calculation of the entanglement entropy, which accounts for the complete microstate structure of the light-cone wave function, might yield different saturation characteristics. Recent works [13, 12] suggest that entanglement entropy could offer a more direct probe of the information storage capacity of the proton.
In contrast, once the same analysis chain is applied to nuclear targets (fixed coupling, identical numerical settings otherwise), the nuclear curves rise faster. The resulting nuclear entropy $S_{A}(x)$ approaches and, within the uncertainties of our inputs, can reach the benchmark line $1/\alpha_{s}$ in a small-$x$ window where the proton stays below. This qualitative separation is stable against reasonable variations of the $Q_{s}$ extraction threshold, geometric inputs, and numerical smoothing; these choices primarily shift the apparent crossing point but do not invert the ordering. Encouragingly, some studies have already explored the correspondence between black hole saturons and the CGC framework [12]. Portions of those discussions align conceptually with the present work, and notably, they propose a role for Goldstone modes arising from the breaking of Poincaré symmetry. The breaking of Poincaré invariance in this context fundamentally originates from the presence of a hard scale, which characterizes the transition from a dilute partonic gas with a spectral gap to an over-occupied, highly saturated classical field. Partons with momenta below this scale become screened, and high-twist effects become significant. This screening of partonic modes disrupts the translational invariance of the partonic gas, analogous to the breaking of Poincaré symmetry in black holes, which gives rise to a Goldstone scale characterized by $f=m_{\text{Planck}}$ with Planck mass $m_{\text{Planck}}$ [3]. However, in the case of QCD, the precise form and interpretation of such a Goldstone scale remain open to further investigation.
@@
In this work, we explored whether a high-energy proton in the small-$x$ regime could qualify as a saturon—a highly occupied field configuration that maximizes entropy under unitarity constraints, in analogy with black holes. We derived the gluon occupation number $N_{g}(x)$ and a thermodynamic entropy $S(x)$ based on analytical and numerical solutions to the BK equation (with both fixed and running coupling), supplemented with an Unruh-inspired temperature $T=Q_{s}/(2\pi)$ and an emergent gluon mass $M_{g}\sim Q_{s}$. Our findings indicate that although both $N_{g}$ and $S$ increase as $x$ decreases, even down to $x\sim 10^{-7}$, they remain significantly below the theoretical saturon bound $1/\alpha_{s}$. This suggests that while the proton exhibits nonlinear saturation dynamics, under the current theoretical framework it does not reach the critical maximal occupancy characteristic of a saturon.
After extending the same framework to nuclear targets at fixed coupling, the nuclear entropy $S_{A}(x)$ rises more sharply and, in our numerical study, can reach the $1/\alpha_{s}$ benchmark in a region where the proton does not. This proton–nucleus contrast identifies nuclei as more favorable environments to probe saturon-like behavior in QCD. Moreover, exploring the connection between the proton and black holes from the perspective of entanglement entropy offers a promising avenue [13].
To bridge the gap between theoretical saturon limits and experimental observables, we identify three promising phenomenological avenues:
1. Entropy-Multiplicity Duality: Theoretical works suggest a link between the entanglement entropy of the initial state and the final-state hadron multiplicity ($S_{ent}\sim\ln\langle N_{ch}\rangle$) [78, 83]. If a QCD system reaches the saturon limit (maximal entropy), particle production should be constrained by this information bound. Precision measurements of charged particle multiplicity distributions $P(N_{ch})$—specifically deviations from Koba-Nielsen-Olesen (KNO) scaling [98] at high multiplicities—could serve as a signature of this entropy saturation.
2. Thermal-like Spectra and Initial-State Correlations: A defining feature of the saturon is its thermodynamic behavior, manifesting as exponential transverse momentum spectra with an effective temperature $T\sim Q_{s}/2\pi$. In high-multiplicity events, observing such “thermal” spectra in the absence of a macroscopic heat bath would support the saturon interpretation. Furthermore, while standard thermalization implies isotropy, the high-occupancy saturon state preserves intrinsic quantum correlations. Specifically, azimuthal correlations ($v_{n}$) are predicted to arise directly from Bose enhancement in the projectile wave function [99], rather than from final-state hydrodynamics. The simultaneous observation of thermal-like spectra (entropy saturation) and these specific initial-state correlations (quantum coherence) would constitute a unique signature of the saturon regime.
3. Unitarity Limit in Diffraction at the EIC: Since our results indicate that nuclei approach the $1/\alpha_{s}$ bound more readily than protons, the upcoming Electron-Ion Collider (EIC) is the ideal discovery machine [100, 101, 102, 103]. However, mere enhancement of diffraction is not unique to the saturon. We anticipate that if the true saturon limit is realized, observables such as the ratio of diffractive to total cross-sections ($\sigma_{diff}/\sigma_{tot}$) should not just increase, but exhibit universal scaling behaviors converging toward the black disk limit ($\sigma_{diff}/\sigma_{tot}\to 1/2$) [104, 105]. Reaching this fundamental unitarity bound would be the definitive signature of the maximal entropy state, distinct from pre-saturation effects.
@@ Acknowledgments
This work has been supported by the National Natural Science Foundation of China (Grant No. 12547118), the Research Program of State Key Laboratory of Heavy Ion Science and Technology, Institute of Modern Physics, Chinese Academy of Sciences (Grant No. HIST2025CS08), and the National Key R&\&D Program of China (Grant No. 2024YFE0109800 and 2024YFE0109802).